# Replacing Softmax Similarity with a Sharpened Angular Similarity: Theory and Practice of Scaling to Billion-Context Attention

### v2

@@ Abstract
Softmax Attention has a quadratic time complexity in sequence length, which becomes prohibitive to run at long contexts, even with highly optimized GPU kernels. For example, FlashAttention (an exact, GPU-optimized implementation of Softmax Attention) cannot complete a single forward–backward pass of a multi-head attention layer once the context exceeds $\sim\!4$ million tokens on an NVIDIA GH200 (96 GB). We introduce RACE Attention, a kernel-inspired alternative to Softmax Attention that is linear in sequence length and embedding dimension. RACE Attention replaces the exponential kernel with a sharpened angular (cosine) similarity, and approximates attention outputs via randomized projections and soft Locality-Sensitive Hashing (LSH). Across language modeling, masked language modeling, and text/image classification, RACE Attention matches or outperforms strong baselines while reducing wall-clock time and memory. In a controlled scale test, it processes up to 12 million tokens during a single forward-backward pass on an NVIDIA GH200 GPU and 75 million tokens on an Intel Xeon® Gold 5220R CPU—well beyond the practical limits of the current state-of-the-art attention implementations. RACE Attention thus offers a practical, theoretically grounded mechanism for outrageously long context windows on today’s hardware. We hope that it gets adopted in practice. Code is open-sourced at https://github.com/sahiljoshi515/RACE_Attention.
@@ 1 Introduction
The Transformer [31, 11] is the backbone of modern sequence modeling across language, vision [23], and speech [19]. We have seen remarkable improvements over the past few years in reasoning and understanding capabilities. Most of these are attributed to the increased parameters of the transformers along with the capability to process longer context windows than before. All this progress, however, rests on a computationally expensive primitive: Softmax Attention, whose time scales quadratically with context length. As models and contexts grow—from multi-document reasoning to long-form code, audio, and video—this quadratic barrier increasingly dictates who can train and deploy capable systems. Industrial labs mitigate the cost with large-scale distributed hardware; most practitioners cannot. There is a growing need for attention mechanisms that are accurate, fast, and memory-efficient. To highlight the limits of Softmax Attention: even with FlashAttention [10]—the state-of-the-art GPU implementation—a single forward–backward pass of a multi-head attention layer (1 batch, 4 heads, embedding dimension of 128) remains computationally and memory intensive and cannot process sequences beyond $\sim 4$ million tokens on an NVIDIA GH200 (96 GB) GPU. Clearly, to achieve an outrageously long context where the target context size is hundreds of millions of tokens or beyond, fundamental rethinking of attention will be required [2].
@@
Linearized and Low-Rank Approximations to Quadratic Attention: Due to the significance of the problem, a very large body of work attempts to accelerate attention by approximating softmax with linear approximations or clever kernel feature maps [15, 6, 24, 25]. Two notable lines of work in the direction of linearly approximating attention are Linear Attention [15] and Performer [6]. Linear Attention replaces the softmax similarity with a simple positive kernel via a feature map, e.g. $\phi(x)=elu(x)+1$. This lets the attention be re-ordered into associative sums, achieving linear computation. Although such a kernel trick reduces computational complexity, it often degrades accuracy, as clearly demonstrated by our experiments in Section 4.1. Performer takes a different approach and cleverly leverages the classical idea of approximating the exponential of an inner product using Random Fourier Features [26]. However, this strategy comes with its own drawbacks. In particular, the method incurs a time complexity that is quadratic in the embedding dimensionality, which offsets many of the intended computational savings. Furthermore, it is well established [1] that approximations based on Random Fourier Features require high-dimensional representations to achieve satisfactory accuracy. Our experimental results in Section 4.2 reinforce this limitation by showing that these methods exhibit poor scalability in practice.
@@
Another category of work replaces the full $N\times N$ attention matrix with a low-rank surrogate. Some methods learn length-wise projections for keys/values (e.g., Linformer), while others use Nyström landmarks to approximate Softmax Attention matrix with a rank-$k$ decomposition (e.g., Nyströmformer). These approaches reduce the leading cost from $\mathcal{O}(N^{2}d)$ to $\mathcal{O}(Nkd)$, at the cost of choosing (and occasionally increasing) $k$ to maintain accuracy. [32, 34]. Moreover, these methods provide no support for autoregressive tasks. As shown in Section 4.1, our method outperforms Linformer in accuracy despite Linformer having $13\%$ more parameters than the other methods. Beyond the empirical shortcomings, a deeper conceptual issue persists: existing approximation approaches lack a rigorous mathematical framework to characterize the trade-offs between efficiency and accuracy. For example, while Performer provides strong kernel-approximation guarantees, a general framework connecting efficiency knobs (e.g., feature count $m$) to downstream accuracy remains limited, and strong accuracy frequently entails large $m$ in practice. As a result, design decisions often appear ad hoc and fragile, leaving methods vulnerable to instability between tasks and settings. Taken together, these limitations explain why, despite the abundance of approximations, Softmax Attention continues to remain the most widely adopted and reliable formulation.
@@
Sparsity is Complementary: We note that there is also a popular line of work [3, 36, 16, 14] that exploits structural information in natural language, with sparsity in attention being among the most widely studied. These approaches are complementary to our proposal, which focus on making the fundamental attention mechanism itself more efficient and mathematically grounded. In principle, our method can be integrated with structural priors such as sparsity to further improve scalability and accuracy. However, since our objective in this paper is to develop fundamentally efficient attention, we will not discuss this line of structural approaches further, instead we view combining them with our method as an important direction for future work.
@@
Our Finding: Standard attention relies on the well-known softmax function, computing
@@
where the softmax is applied row-wise so that attention weights are nonnegative and sum to one. In this paper, we propose a surprising alternative to the softmax—namely, a high-degree monomial of an angular kernel based on cosine geometry:
@@
Eq. (2) should be read as an informal analogue to Eq. (1), where the angular kernel replaces the exponential. A more precise definition, with explicit cosine normalization and row-wise normalization, is given in Section 3.1. We argue that, for sufficiently large values of $\gamma$, this formulation closely mimics the behavior of softmax and refer to it as Angular Attention in the subsequent sections. Importantly, it admits a linear-time approximation algorithm. In particular, we leverage the connection (Section 2.2) between Repeated Arrays-of-Count Estimators (RACE) [8, 18] and the angular kernel to design our algorithm in Section 3.2. We therefore refer to our proposed method as RACE Attention.
RACE Attention is a drop-in replacement for Softmax Attention. We evaluate it in Transformers on language modeling, masked language modeling, and text/image classification (Section 4.1). By reframing similarity around an angular kernel and using differentiable LSH-based sketches, it provides a principled alternative that supports very long contexts on commodity hardware. The sketching view keeps constant factors small: each query mixes with only a fixed bank of $S=LR$ bucket summaries rather than all $N$ keys. Since we never materialize the full attention matrix, the working set stays compact and activation memory drops, enabling much longer sequences with reduced latency. In contrast, Softmax Attention retains full $Q,K,V,O$ tensors, preventing processing of much longer sequences at comparable speed (Section 4.2). In addition to our novel findings about RACE Attention and rigorous supporting experimental evidence, we provide the following:
  • Long-context scaling: We demonstrate scaling at context lengths far beyond prior attention mechanisms—up to 75 million tokens on a CPU and 12 million tokens on a GPU. To the best of our knowledge, this regime has not been reached by existing attention variants on a single device.
  • Trainable RACE: We make the sketch differentiable by using soft assignments to the hypercube corners (rather than hard hashes), enabling end-to-end learning.
  • CPU/GPU pre-training: We support both causal (autoregressive) and non-causal (bidirectional) pre-training on CPU and GPU. For CPU workloads, we provide a custom OpenMP kernel that computes the causal prefix operations in a single pass using efficient streaming algorithm, enabling linear-time, memory-efficient training.
  • Theory and ablations: We provide approximation guarantees in Section 3.3 inherited from LSH and analyze how sketch parameters—number of hash tables $L$, buckets per table RR—govern variance–accuracy trade-offs. Comprehensive ablations quantify accuracy, speed, and memory as these knobs are varied.
@@ 2.1 Locality-Sensitive Hashing (LSH)
$\begin{cases}\mathrm{Sim}(x,y)\geq S_{0}\;\Rightarrow\;\Pr_{h\sim\mathcal{H}}[h(x)=h(y)]\geq p_{1},\\[2.0pt] \mathrm{Sim}(x,y)\leq cS_{0}\;\Rightarrow\;\Pr_{h\sim\mathcal{H}}[h(x)=h(y)]\leq p_{2},\end{cases}\quad p_{1}>p_{2},\;c<1.$
Such families enable sublinear-time approximate nearest-neighbor data structures. A convenient sufficient condition—satisfied by SimHash and WTA [4, 35, 5]—is that the collision probability is a monotone function of similarity, $\Pr_{h\sim\mathcal{H}}[h(x)=h(y)]=f(\mathrm{Sim}(x,y))$ with $f$ increasing.
Figure 1: This figure demonstrates the difference between the linear complexity of RACE Attention and the quadratic complexity of Softmax Attention mechanism. Specifically, we highlight how the final representation $o_{5}$ is computed under Softmax versus RACE. In Softmax, the entire fifth column of the attention score matrix is required. In contrast, RACE does not require the full matrix; instead, it aggregates statistics within LSH-mapped buckets, utilizing the appropriate collision probability $\alpha$ to compute $o_{5}$.
@@ 3 Introducing RACE Attention
3.1 Softmax-Like Similarities that Admit Linear-Time Estimation
Given a sequence of $N$ tokens, a Transformer produces for each position $i$ a query $Q_{i}\in\mathbb{R}^{d}$, and for every position $j$ a key $K_{j}\in\mathbb{R}^{d}$ and a value $V_{j}\in\mathbb{R}^{d}$. The output at position $i$ is a weighted average of the values, where the weight on $V_{j}$ reflects the relevance of token $j$ to token $i$. In the standard formulation [31], relevance is computed via the scaled dot product given by Eq. (1). This choice guarantees two useful properties of the attention weights: (i) non-negativity and (ii) they sum to one, so $O_{i}$ is a convex combination of the values. Equally important, the exponential introduces a strong non-linear mapping from similarity scores to attention weights, amplifying small score differences. This observation suggests a broader view: attention weights can be derived from any normalized highly non-linear (exponential like) similarity function. Let $\mathrm{sim}:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$ be any non-negative similarity function. We can define normalized similarity attention as
Our quest is for finding non-linear (softmax-like) similarity kernels that admit accurate linear-time estimation, eliminating the quadratic cost of attention in both training and inference. We argue that a good starting point is a well known LSHable [8, 7] angular similarity. It is well behaved and normalized, in particular, it depends only on the angle between the vectors $Q_{i}$ and $K_{j}$ and is invariant to their norms: $\mathrm{sim}(Q_{i},K_{j})=1-\cos^{-1}\!\left(\frac{Q_{i}^{\top}K_{j}}{\lVert Q_{i}\rVert\,\lVert K_{j}\rVert}\right)/\pi$
However, unlike exponential in softmax, the raw angular kernel is relatively flat near high similarity values, reducing its ability to sharply discriminate between nearly aligned vectors. To increase contrast, we propose to exponentiate the angular kernel with a sharpening parameter $\gamma$, which accentuates differences among highly similar pairs. After sharpening the kernel, the similarity function becomes as follows:
In Figure 2, we show that for sufficiently large $\gamma$, the angular kernel becomes almost indistinguishable from softmax kernel. This is expected because a high degree monomial like $x^{12}$ behaves similarly to an exponential.
@@ 3.1 Softmax-Like Similarities that Admit Linear-Time Estimation
In its current form, evaluating the attention with similarity given by Eq. (4) is no different from softmax. It naively still requires all $N^{2}$ query–key interactions. Fortunately, any constant exponentiation of angular kernel, belongs to a family, that admits efficient kernel density estimation using RACE sketches [8]. In particular, we will use LSH-based RACE sketches to approximate the kernel in linear time obtaining an algorithmically efficient alternative to Softmax Attention!
Figure 2: Comparison of Softmax and Angular kernels at different sharpening levels $\gamma$. As $\gamma$ (or non-linearity) increases, Angular transitions from flat similarity scores to a sharper distribution, recovering behavior similar to the exponential in the Softmax.

### v3

# RACE Attention: A Strictly Linear-Time Attention for Long-Sequence Training

@@ Abstract
Softmax Attention has a quadratic time complexity in sequence length, which becomes prohibitive to run at long contexts, even with highly optimized GPU kernels. For example, FlashAttention-2/3 (exact, GPU-optimized implementations of Softmax Attention) cannot complete a single forward–backward pass of a single attention layer once the context exceeds $\sim\!4$ million tokens on an NVIDIA GH200 (96 GB). We introduce Repeated Arrays-of-Count Estimators (RACE) Attention, a kernel-inspired alternative to Softmax Attention that is strictly linear in sequence length and embedding size. RACE Attention replaces the exponential kernel with a sharpened angular similarity, and approximates attention outputs via Gaussian random projections and soft Locality-Sensitive Hashing (LSH), avoiding construction of the full attention matrix. Across language modeling, masked language modeling, and text/image classification, RACE Attention matches or outperforms strong baselines up to 6464K seqeuence length while reducing wall-clock time and memory usage. In addition, we conduct a controlled scaling study on a single attention layer and demonstrate processing of up to 12 million tokens on an NVIDIA GH200 GPU and 75 million tokens on an Intel Xeon® Gold 5220R CPU in a single forward–backward pass, which is well beyond the capabilities of current state-of-the-art attention implementations. RACE Attention thus offers a practical and theoretically grounded mechanism for long-context training on today’s hardware. We release our code at https://github.com/sahiljoshi515/RACE_Attention.
@@ 1 Introduction
The Transformer [37, 13] is the backbone of modern sequence modeling across language, vision [26], and speech [22]. We have seen remarkable improvements over the past few years in reasoning and understanding capabilities. Most of these are attributed to the increased parameters of the transformers along with the capability to process longer context windows than before. All this progress, however, rests on a computationally expensive primitive: Softmax Attention, whose time scales quadratically with context length. As models and contexts grow, spanning multi-document reasoning, long-form code, audio, and video, the quadratic barrier increasingly dictates who can train and deploy capable systems. Industrial labs mitigate the cost with large-scale distributed hardware; most practitioners cannot. There is a growing need for attention mechanisms that are accurate, fast, and memory-efficient. To highlight the limits of Softmax Attention: even with FlashAttention-2/3 [12, 33], the state-of-the-art GPU implementations, a single forward–backward pass of a single attention layer (1 batch, 4 heads, 128 embedding size) remains computationally and memory intensive and cannot process sequences beyond $\sim 4$ million tokens on an NVIDIA GH200 (96 GB) GPU. Clearly, to achieve an outrageously long context where the target context size is hundreds of millions of tokens or beyond, fundamental rethinking of attention [2] will be required.
Figure 1: This figure demonstrates the difference between the linear complexity of RACE Attention and the quadratic complexity of Softmax Attention. Specifically, we highlight how the final representation $o_{5}$ is computed under Softmax versus RACE. In Softmax, the entire fifth column of the attention score matrix is required. In contrast, RACE does not require the full matrix; instead, it aggregates statistics within LSH-mapped buckets, utilizing the appropriate collision probability $\alpha$ to compute $o_{5}$.
@@
Linearized and Low-Rank Approximations to Quadratic Attention: Due to the significance of the problem, a very large body of work attempts to accelerate attention by approximating softmax with linear approximations [43] or clever kernel feature maps [18, 7, 27, 29]. A work more closely related to ours is YOSO Attention [43], which also approximates a powered angular kernel using hard LSH. Despite this same kernel formulation, the underlying estimators differ substantially. RACE Attention leverages a smooth, differentiable relaxation of classical RACE-based hashing [10] to approximate the powered angular kernel. More importantly, YOSO lacks formal theoretical guarantees on its approximation quality and a mechanism that enables causal language modeling. RACE addresses both of these limitations: Sections 3.2 and 3.3 develop a theoretically grounded estimator with quantifiable approximation error, and Algorithm 2 presents a practical attention mechanism that naturally supports causal settings. We defer an in depth comparison between RACE and YOSO to Section 3.1. Two notable lines of work in the direction of linearly approximating attention are Linear Attention [18] and Performer [7]. Linear Attention replaces the softmax similarity with a simple positive kernel via a feature map, e.g. $\phi(x)=elu(x)+1$. This lets the attention be re-ordered into associative sums, achieving linear computation. Although such a kernel trick reduces computational complexity, it often degrades accuracy, as clearly demonstrated by our experiments in Section 4. Performer takes a different approach and cleverly leverages the classical idea of approximating the exponential of an inner product using Random Fourier Features [30]. However, this strategy comes with its own drawbacks. In particular, the method incurs a time complexity that is quadratic in the embedding size, which offsets many of the intended computational savings. Furthermore, it is well established that approximations based on Random Fourier Features require high-dimensional representations to achieve satisfactory accuracy [1]. Our experimental results in fig. 4 reinforce this limitation by showing that these methods exhibit poor scalability in practice.
@@
Another category of work replaces the full $N\times N$ attention matrix with a low-rank surrogate. Some methods learn length-wise projections for keys/values (e.g., Linformer), while others use Nyström landmarks to approximate Softmax Attention matrix with a rank-$k$ decomposition (e.g., Nyströmformer). These approaches reduce the leading cost from $O(N^{2}d)$ to $O(Nkd)$, at the cost of choosing (and occasionally increasing) $k$ to maintain accuracy [38, 40]. Moreover, these methods provide no support for autoregressive tasks. As shown in Section 4, our method outperforms Linformer in accuracy despite Linformer having $13\%$ more parameters than the other methods. Beyond the empirical shortcomings, a deeper conceptual issue persists: existing approximation approaches lack a rigorous mathematical framework to characterize the trade-offs between efficiency and accuracy. For example, while Performer provides strong kernel-approximation guarantees, a general framework connecting efficiency knobs (e.g., feature count $m$) to downstream accuracy remains limited, and strong accuracy frequently entails large $m$ in practice. As a result, design decisions often appear ad hoc and fragile, leaving methods vulnerable to instability between tasks and settings. Taken together, these limitations explain why, despite the abundance of approximations, Softmax Attention continues to remain the most widely adopted and reliable formulation.
@@
Sparsity is Complementary: We note that there is also a popular line of work [3, 42, 19, 16, 28] that exploits structural information in natural language, with sparsity in attention being among the most widely studied. These approaches are complementary to our proposal, which focus on making the attention mechanism itself more efficient and mathematically grounded. In principle, our method can be integrated with structural priors such as sparsity to further improve scalability and accuracy. However, since our objective in this paper is to develop fundamentally efficient attention, we will not discuss this line of structural approaches further, instead we view combining them with our method as an important direction for future work.
@@
Key Idea: Standard attention relies on the well-known softmax function, computing
@@
where the softmax is applied row-wise so that attention weights are nonnegative and sum to one. In this paper, we leverage an alternative to the softmax—namely, a higher-degree monomial of an angular kernel based on cosine geometry:
@@
Eq. 2 should be read as an informal analogue to Eq. 1, where the angular kernel replaces the exponential. A more precise definition, with explicit cosine normalization and row-wise normalization, is given in Section 3.1. We argue that, for sufficiently large values of $\gamma$, this formulation closely mimics the behavior of softmax and refer to it as Angular Attention in the subsequent sections. Importantly, it admits a linear-time approximation algorithm. In particular, we leverage the connection (Section 2.2) between Repeated Arrays-of-Count Estimators (RACE) [10, 21] and the angular kernel to design our algorithm in Section 3.2. We therefore refer to our proposed method as RACE Attention.
RACE Attention is a drop-in replacement for Softmax Attention. We evaluate it in Transformers on causal language modeling, masked language modeling, and text/image classification (Section 4 and Appendix 5.1). By reframing similarity around a powered angular kernel and using differentiable LSH-based sketches, it provides a principled alternative that supports very long contexts on commodity hardware. The sketching view keeps constant factors small: each query mixes with only a fixed bank of $S=LR$ bucket summaries rather than all $N$ keys. Since we never materialize the full attention matrix, the working set stays compact and activation memory drops, enabling much longer sequences with reduced latency. In contrast, FlashAttention-2/3 [12, 33] reduce the memory footprint of attention via tiling, but still require computing all key-query interactions, preventing processing of much longer sequences at comparable speed, as shown in fig. 4. In addition to our novel findings about RACE Attention and rigorous supporting experimental evidence, we provide the following:
I. Long-context scaling: RACE scales to sequence lengths far beyond prior attention mechanisms, processing up to 75M tokens on CPU and 12M tokens on GPU in a single forward–backward pass of a single attention layer, using the same hyperparameters as in our accuracy evaluations.
II. Trainable RACE: We introduce a differentiable sketch by replacing hard hashing with smooth soft assignments over hypercube corners, enabling end-to-end training.
III. CPU/GPU pre-training: Our approach supports both causal (autoregressive) and non-causal (bidirectional) pre-training on CPU and GPU, with custom OpenMP/CUDA kernels that compute causal prefix operations in a single streaming pass for linear-time, memory-efficient training.
IV. Theoretical Insights: Section 3.3 establishes approximation guarantees based on the LSH framework and analyzes how the parameters $L$ (number of hash tables) and $R$ (buckets per table) govern the variance–accuracy tradeoff.
@@ 2.1 Locality-Sensitive Hashing (LSH)
$\begin{cases}\mathrm{Sim}(x,y)\geq S_{0}\;\Rightarrow\;\Pr_{h\sim\mathcal{H}}[h(x)=h(y)]\geq p_{1},\\[2.0pt] \mathrm{Sim}(x,y)\leq cS_{0}\;\Rightarrow\;\Pr_{h\sim\mathcal{H}}[h(x)=h(y)]\leq p_{2},\end{cases}$
where $p_{1}>p_{2}$ and $c<1$. Such families enable sublinear-time approximate nearest-neighbor data structures. A convenient sufficient condition, satisfied by SimHash and WTA [5, 41, 6], is that the collision probability is a monotone function of similarity, $\Pr_{h\sim\mathcal{H}}[h(x)=h(y)]=f(\mathrm{Sim}(x,y))$ with $f$ increasing.
@@ 3 Introducing RACE Attention
$\phi^{(\ell)}(x)]_{r}=\frac{\exp\{\beta\,(\tanh(W^{(\ell)}x))^{\top}v_{r}\}}{\sum_{r^{\prime}}\exp\{\beta\,(\tanh(W^{(\ell)}x))^{\top}v_{r^{\prime}}\}},\quad x\in\{Q_{i},K_{j}\}$
@@
3.1 Softmax-Like Similarities that Admit Linear-Time Estimation
Figure 2: Frobenius error between Angular and Softmax Attention vs. $\gamma$.
Given a sequence of $N$ tokens, a Transformer produces for each position $i$ a query $Q_{i}\in\mathbb{R}^{d}$, and for every position $j$ a key $K_{j}\in\mathbb{R}^{d}$ and a value $V_{j}\in\mathbb{R}^{d}$. The output at position $i$ is a weighted sum of the values, where the weight on $V_{j}$ reflects the relevance of token $j$ to token $i$. In the standard formulation [37], relevance is computed via the scaled dot product given by Eq. 1. This choice guarantees two useful properties of the attention weights: (i) non-negativity and (ii) they sum to one, so $O_{i}$ is a convex combination of the values. Equally important, the exponential introduces a strong non-linear mapping from similarity scores to attention weights, amplifying small score differences. This observation suggests a broader view: attention weights can be derived from any normalized highly non-linear (exponential like) similarity function. Let $\mathrm{sim}:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$ be any non-negative similarity function. We can define normalized similarity attention as
Our quest is for finding non-linear (softmax-like) similarity kernels that admit accurate linear-time estimation, eliminating the quadratic cost of attention in both training and inference. Similar to YOSO [43], we argue that a good starting point is a well known LSHable [10, 8] angular similarity. It is well behaved and normalized, in particular, it depends only on the angle between the vectors $Q_{i}$ and $K_{j}$ and is invariant to their norms: $\mathrm{sim}(Q_{i},K_{j})=1-\cos^{-1}\!\left(\frac{Q_{i}^{\top}K_{j}}{\lVert Q_{i}\rVert\,\lVert K_{j}\rVert}\right)/\pi$. However, unlike exponential in softmax, the raw angular kernel is relatively flat near high similarity values, reducing its ability to sharply discriminate between nearly aligned vectors. To increase contrast, we propose to exponentiate the angular kernel with a sharpening parameter $\gamma$, which accentuates differences among highly similar pairs. After sharpening the kernel, the similarity function becomes as follows:
In fig. 3, we show that for sufficiently large $\gamma$, the angular kernel becomes almost indistinguishable from softmax kernel. This is expected because a higher degree monomial like $x^{12}$ behaves similarly to an exponential. Furthermore, in fig. 2 we plot the frobenius error between Angular and Softmax Attention. The error sharply decreases as $\gamma$ increases, demonstrating that softmax-level sharpness can be achieved with modest polynomial degree (e.g., $\gamma=8$).
Figure 3: Comparison of Softmax and Angular kernels at different sharpening levels $\gamma$. As $\gamma$ (non-linearity) increases, Angular transitions from flat similarity scores to a sharper distribution, recovering behavior similar to the exponential in Softmax.
In its current form, evaluating the attention with similarity given by Eq. 4 is no different from softmax. It naively still requires all $N^{2}$ query-key interactions. Fortunately, any constant exponentiation of angular kernel, belongs to a family, that admits efficient kernel density estimation using RACE sketches [10], and we use these sketches to approximate the kernel in linear time obtaining an algorithmically efficient alternative to Softmax Attention!
Finally, as outlined in Section 1, YOSO [43] and our method operate on the same kernel in Eq. 4, but differ fundamentally in how attention is computed. YOSO employs hard LSH to generate Bernoulli collision indicators, yielding an unbiased estimator of the unnormalized attention numerator $\sum_{j=1}^{N}\mathrm{sim}(Q_{i},K_{j})V_{j}$ in Eq. 3. It then applies a post-hoc $\ell_{2}$ normalization, which does not correspond to standard attention normalization. Since this hard formulation is non-differentiable, YOSO relies on surrogate lower-bound gradients derived from additional Bernoulli samples rather than directly differentiating the kernel. This introduces quadratic complexity in the embedding dimension $d$, resulting in poor scalability during end-to-end training (see fig. 5(c)). To see how RACE constructs a smooth, differentiable LSH-based estimator that is linear in $d$ in contrast to YOSO, refer to Section 3.2.